Shareholders in firm B are paid the market value of their firm in shares of stock from firm A. The earnings of the combined firm are $68,000. Information on each firm, prior to merger is as follows:

Firm A Firm B
Number of outstanding shares 30000 22000
Price per share $32.00 $25.00
Debt $0 $0
Total earnings $36,000.00 $30,000.00



What is the net present value of acquiring firm B in an all stock merger?

Multiple Choice

  • $6,375

  • $17,188

  • $9,533

  • $13,221

  • $10,000
Answer #1

value of synergy = $ 2,000 ( 68000-36000-30000)

EPS

firm A = 36000 / 30000 = $ 1.2

firm B = 30000 / 22000 = $ 1.36

PE ratio

firm A = 32 / 1.2 = 26.6

firm B = 25 / 1.36 = 18.38

multiple of earnings to price

firm A= 100/26.6 = 3.76

firm B= 100/ 18.3 = 5.46

assume synergies value is $ 10,000 (2,000 * 5)

Value A = 32 * 30000 = $ 960,000

Value B = 25 * 22000 = $ 550,000

V B * = $ 550,000 + 10,000 = $ 560,000

A must sell stock at the current market price of $ 32 . Firm A needs to issue

550,000 / 32 = 17,188 shares

after the merger there will be 30,000 + 17, 188 = 47,188 shares

price per share will be

( 960,000+560,000+10,000)/47,188 = 32.21

price paid = 560,000 - 17,188 * 32.21 = $ 6374.5 or $ 6375.